OpenStudy (anonymous):

Show steps? (x - 3) /x = (x - 3) / (x - 6)

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

\[(x-3)/x = (x-3)/(x-6) \rightarrow x = 3\]

OpenStudy (anonymous):

are u trying to solve for x?

OpenStudy (anonymous):

Yes. I'm trying to solve for x. Can you show me the steps on how you got x=3?

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

your welcome ;0